# F-91191 Gif-sur-Yvette cedex, France

### v1

@@ 4.2 Coulomb region
Consequently, for $d=4$, we see that this region can depart from the expected peeling property only at tree level, $L=0$, and moreover that it can occur only for the NP scalar $\Psi_{0}$, which is expected to scale as $|\vec{x}|^{-5}$ if the peeling property holds. We note that this argument does not cover logarithmic tail contributions, arising from the interplay of $d$-dependent exponent in Eq. (70) and IR divergences from higher-order contributions. We will see later that, nevertheless, the tail terms do not contribute in the Coulomb-U-SE region.
@@ 4.2.1 Coulomb region: Leading-soft contribution
This is effectively an expansion around the null direction connecting the scattering process and the observer, per Eq. (40).
@@
$\begin{split}\int\!d^{d}ke^{-ik\cdot x}f_{R}(k^{\mu})\Big|_{\text{Coulomb}}&=\int\!dn\int_{-\infty}^{+\infty}\!d\tilde{\omega}\int_{0}^{\infty}d\mathsf{k}\,\mathsf{k}^{d-1+\alpha}e^{-i\mathsf{k}|\vec{x}|(\tilde{\omega}-n\cdot r)}\tilde{f}_{R}(\tilde{\omega},n)\\ &=\frac{\Gamma\left[d+\alpha\right]}{(ir)^{d+\alpha}}\int\!dn\int_{-\infty}^{+\infty}\!d\tilde{\omega}\,\frac{\tilde{f}_{R}(\tilde{\omega},n)}{(\tilde{\omega}-n\cdot r)^{d+\alpha}}=0\ ,\end{split}$ (76)
where setting $t=|\vec{x}|$ in the first equality is a consequence of the restriction to the Coulomb region. The last equality follows from Cauchy’s theorem together with integration contour on the upper-half plane. The singularity on the real axis can be made integrable by a careful choice of the dimensional regularization parameter $d$.
@@ 4.2.2 Coulomb region: subleading-soft and peeling violation
We argued that, because gravitons in the Coulomb region have momenta $O(1/|\vec{x}|)$, their contribution to the metric is governed by the soft expansion of the source. Having discussed the leading-order contributions in the previous section, we now proceed to discuss the first subleading contributions, which are more subtle. Our focus will be the terms that are nonanalytic in the components of the outgoing graviton momentum near the origin. We will find agreement with the results of Refs. Christodoulou:2002 .
Similarly to the leading soft terms, the form of the subleading soft terms are universal Saha:2019tub , independent of the details of the scattering process, when expressed in terms of the incoming and outgoing momenta: [^18] The form of the subleading soft theorem may not look covariant. On the other hand, all naively covariant deformations are all equivalent to this order in the soft expansion, e.g. $\log k\cdot Q\sim\log\omega+\mathcal{O}(\omega^{0})$.
@@
and $Q_{3}^{\prime\mu}=K^{\mu}$ is the total radiated energy during the scattering and the dots stand for the leading soft contributions. These terms originate both from the off-shell momentum-space source as well as from the Fourier-transform to impact-parameter space, from the ultra-soft-exchange region. [^19] On the last line we introduced $\ln(\omega+i\epsilon)\rightarrow\frac{1}{2}\ln(k^{2}+i\epsilon)$ because of the proportionality with the leading order soft factor and because of the consistency of the coefficients with Weinberg’s exponentiation of IR divergences. We also note that the $i\epsilon$ deformations of singularities are such that the integration contour is not pinched.
The soft theorem in Eq. (77) has been derived in $d=4$. This is a clear obstruction to performing this computation within dimensional regularization and the method of regions. Thus, in this case, we are forced to perform to compute the contribution of the Coulomb soft region using cut-off regularization. Power counting suggests that such integrals are UV power-divergent as the cutoff is taken $\lambda\rightarrow+\infty$. As discussed below Eq. (34), such divergences should cancel when the Coulomb and radiation regions are added up, leaving finite terms that depend on the parameters of the integral. Furthermore, we will ignore the retarded contributions (i.e. with denominators $(Q_{i}\cdot k+i\epsilon)$ and $\log(\omega+i\epsilon)$) because they vanish in dimensional regularization, see discussion around Eq. (76). Thus, in other regularizations, we expect them to be UV sensitive and cancel completely against counterparts in the radiation region.
@@
for the Weyl tensor, where the restriction to $t=|\vec{x}|$ is a consequence of expanding in the Coulomb region. We notice that, by computing the $\omega$ integration by Cauchy’s theorem, we are sensitive only to the on-shell poles of the retarded graviton propagator because $t=|\vec{x}|>0$. Thus, all off-shell terms ($\propto k^{2}$) give vanishing contributions. One can show via direct evaluation in dimensional regularization that all the integrals with retarded matter propagators (i.e. with denominators $(Q_{i}\cdot k+i\epsilon)$ and $\log(\omega+i\epsilon)$) give scaleless integrals, see Eq. (76). [^22] As already briefly mentioned below Eq. (78), this statement hides some subtleties. Indeed, by using a different regulator than dimensional regularization, one may find that the retarded contributions give a non-vanishing contribution. Such terms are, however, expected to cancel exactly with analogous regulator-dependent terms in the radiation region. Thus, the details of the outgoing particles – in particular the impulse – do not contribute to the Coulomb region, recovering the conclusion of the scaling argument that the Coulomb (and U-SE) region is tree-level exact.
@@
As briefly mentioned in footnote 20, to enforce integration in the Coulomb region, we relate the cutoff $\lambda$ and $|\vec{x}|$ as $\lambda=\frac{\Lambda}{|\vec{x}|}$ with $\Lambda\gg 1$. The resulting metric is polyhomogeneous, of the form discussed in e.g. Ref. Geiller:2024ryw , with an overall power of $\frac{1}{|\vec{x}|^{2}}$ for the metric and $\frac{1}{|\vec{x}|^{4}}$ for the linearized Weyl tensor, with $\log|\vec{x}|$ enhancements.
To understand which terms may cancel between Eqs. (86) and (86) and their counterparts for $\mathsf{k}>\lambda$, we explore how coefficients of the various tensor structures change in these equations under rescalings and shifts of the cutoff $\lambda$. To this end, we focus on the vector integral contributing to the metric, Eq. (86). Using the relation explained in footnote 20 between $\lambda$ and $|\vec{x}|$, $\lambda=\frac{\Lambda}{|\vec{x}|}$ with $\Lambda\gg 1$ that enforces the restriction to the Coulomb region, it is easy to see that only terms proportional to $L^{\mu}$ can be changed by $\Lambda\rightarrow a\Lambda+b$ with fixed $a$ and $b$. Thus, all terms proportional to $N^{\mu}$ are scheme-independent and should not be expected to be affected by radiation-region contributions that cancel the cutoff dependence. These are the terms contributing to $\Psi_{0}$ and $\Psi_{1}$ NP scalars.
While a similar analysis of the two-tensor integral in Eq. (87) contributing to the Weyl tensor suggests a much stronger scheme dependence. However, since the linearized Weyl tensor is simply given by suitably antisymmetrized derivatives of the metric, its scheme dependence is substantially weaker per our discussion above. In particular, all terms that contain an $N^{\mu}$ coming from the metric are scheme-independent. The $N^{\mu}N^{\nu}$ structure has this property.
@@
the peeling violation of Ref. Damour:1985cm depends on the $A_{ij}$ and that of Ref. Christodoulou:2002 depends on $B_{ij}$. [^23] We would like to thank T. Damour for extensive discussion on this point. We expect that the result of Ref. Damour:1985cm also originates from external gravitons in the Coulomb region. In particular, a dimensional analysis argument paralleling the one outlined at the beginning of this section suggests that a further source of peeling violation could be the subleading term in the soft-exchange region ($\sim\mathsf{k}^{0}$), which we have not considered in our analysis. An alternative origin of the peeling violating terms of Ref. Damour:1985cm could be terms in the Weyl tensor that are nonlinear in the metric fluctuations, in particular terms that depend on the linearized Schwarzschild metric of both incoming particles. We leave these for a future investigation.
@@
Before proceeding to the next section, let us briefly return to the scheme dependence of the contributions to $\Psi_{0}$ and $\Psi_{1}$ computed above. As we already explained, all contributions to the NP scalars coming from the terms proportional to the $N^{\mu}$ in the metric cannot be changed by adjusting the cutoff separating the Coulomb and radiation regions. The disappearance of the cutoff in Eq. (89) reinforces this conclusion. A further indication that $\Psi_{0}$, as computed above, is scheme-independent will be given in the following section, where we will show that the radiation region cannot yield $\Psi_{0}\sim 1/|\vec{x}|^{4}$ at the same order in $G$ as the Coulomb region. $\Psi_{1}$, on the other hand, exhibits explicit cutoff dependence of the schematic form $\Psi_{1}\sim G^{2}/|\vec{x}|^{4}\ln\lambda$. We expect that this dependence will cancel against divergences in the soft integration of the radiation region arising from logarithmic dependence on the retarded time. Moreover, while we kept non-logarithmic terms in Eq. (91), they should not be taken too seriously because they can be changed by altering the choice of cutoff in footnote 20.
@@ 4.3 Radiation regions: the analytic contributions
Having analyzed the contribution of the Coulomb region to the NP scalars and recovered the departure from Sachs’s peeling property found from different considerations in Refs. Damour:1985cm ; Christodoulou:2002 , we now proceed to discuss the contribution of the radiation region. We will show that, in this region, the leading $1/r$ behavior of the Newman–Penrose scalars is determined entirely by the on-shell metric. Moreover, the components of the momentum-space metric that are analytic in the graviton virtuality $k^{2}$ yield position-space contributions that satisfy Sachs’s peeling property. The non-analytic terms will be examined in the next section.
As we discussed, the metric in Fourier space is given by the graviton form factor, with the subtraction of certain discontinuities. The external matter lines are on-shell and amputated through LSZ reduction, while the graviton line is off-shell and not amputated. Since we assume the current $J^{\mu\nu}$ is analytic for the purpose of this section, the integral over $\omega$ can be evaluated using residue theorem. There are contributions from the pole of the external graviton propagator, as well as from possible singularities in $\omega$ within the form factor. The LSZ reduction guarantees that such singularities are at $k^{2}\neq 0$. We will consider them separately and find that the pole of the external propagator contributes to the leading $1/|\vec{x}|$ faloff, while all the others lead to a faster (exponential) falloff.
@@ 4.3.1 The Fourier transform: retarded graviton poles
We begin by considering the pole of the external propagator and use it to evaluate the $\omega$ integral; this effectively sets the source on-shell. We make this choice so that the contour of the remaining $\mathsf{k}$ integral does not pass through any of the remaining poles. Given Eq. (3.1), the result is
@@
Since we are in the radiation region ($k\sim|\vec{x}|^{0}$), we expand ${\widetilde{h}}^{\mu\nu}(k)$, or equivalently, the source $J^{\mu\nu}(k)$, in the angular variable $y$, around y=0y=0.242424The Watson’s lemma guarantees that it is sufficient to consider the asymptotic expansion of the source around such point. A rescaling of $y$ turns this into a series expansion in $|\vec{x}|^{-1}$: each $y$ numerator factor pulls down a factor of $(\mathsf{k}|\vec{x}|)^{-1}$. The expansion in $y$ also generates a polynomial in $n_{\perp}$; the integral over these variables is carried out via identities like Eq. (37). We are left with a non-trivial integral over $\mathsf{k}$.
@@
We can also compute subleading contributions to the metric in the $|\vec{x}|^{-1}$ expansion by keeping higher-order terms in the expansion in $y$. In dimensional regularization at finite $\epsilon=(4-d)/2$, and in the strict on-shell limit enforced by localizing onto the external propagator pole, the momentum-space metric $\tilde{h}^{\mu\nu}(k)$ becomes independent of $k^{2}$. This suggests that we should not expect any departure from the leading-order behavior described above, at least if the initial state contains only two matter particles and no gravitational radiation.
@@ 4.3.2 The Fourier transform: contribution of other poles
Having understood the first and last terms in Eq. (92) and their relation to the KMOC representation of the waveform, let us proceed to the second and third terms in that equation. They depend on the advanced time, which is parametrically large in the limit we are interested in, $v=t\,+|\vec{x}|\sim 2|\vec{x}|\rightarrow\infty$. Thus the $\mathsf{k}$ integral is dominated by $\mathsf{k}\sim{\cal O}(|\vec{x}|^{-1})$ which, together with the on-shell condition imposed by the external graviton pole, implies that
$k^{\mu}\sim{\cal O}(|\vec{x}|^{-1})\ .$ (94)
Thus, these integrals are dominated by graviton momenta in the Coulomb region, which we already discussed in Secs. 4.2.1 and 4.2.2. In the radiation region, where $k^{\mu}\sim{\cal O}(|\vec{x}|^{0})$, their contribution is exponentially suppressed by the Fourier phase factor.
We now discuss the contribution of other possible poles in $\omega$ that originate from the momentum-space metric ${\widetilde{h}}^{\mu\nu}(k)$. All such poles will set $\mathsf{k}=a\omega$ where $a$ is some function of scalar products. This coefficient cannot be unity because because such a value also sets the external graviton on-shell and such a higher-order pole is in disagreement with known properties of 5-point S-matrix elements. The residue of the Fourier-transform integrand at this pole is proportional to $\exp[\pm i\omega(t\pm a|\vec{x}|)]$. The exponent is parametrically large in all terms. Thus, as for the advanced poles, the contribution of these poles to the Fourier transform is from the Coulomb region, $k^{\mu}\sim{\cal O}(|\vec{x}|^{-1})$. It is exponentially suppressed in the radiation region.

### v2

@@ 4.2 Coulomb region
Consequently, for $d=4$, we see that this region can depart from the expected peeling property only at tree level, $L=0$, and moreover that it can occur only for the NP scalar $\Psi_{0}$, which is expected to scale as $|\vec{x}|^{-5}$ if the peeling property holds. We note that this argument does not cover logarithmic contributions. Indeed, we will show that even for $\Psi_{1}$, we will reproduce a known logarithmic departure from the peeling conditions.
@@ 4.2.1 Coulomb region: Leading-soft contribution
$\begin{split}\int\!d^{d}ke^{-ik\cdot x}f_{R}(k^{\mu})\Big|_{\text{Coulomb}}&=\int\!dn\int_{-\infty}^{+\infty}\!d\tilde{\omega}\int_{0}^{\infty}d\mathsf{k}\,\mathsf{k}^{d-1+\alpha}e^{-i\mathsf{k}|\vec{x}|(\tilde{\omega}-n\cdot r)}\tilde{f}_{R}(\tilde{\omega},n)\\ &=\frac{\Gamma\left[d+\alpha\right]}{(i|\vec{x}|)^{d+\alpha}}\int\!dn\int_{-\infty}^{+\infty}\!d\tilde{\omega}\,\frac{\tilde{f}_{R}(\tilde{\omega},n)}{(\tilde{\omega}-n\cdot r)^{d+\alpha}}=0\ ,\end{split}$ (74)
where setting $t=|\vec{x}|$ in the first equality is a consequence of the restriction to the Coulomb region. The singularity on the real axis can be made integrable by a careful choice of the dimensional regularization parameter $d$. The last equality follows from Cauchy’s theorem, by deformation of the integration contour on the upper-half plane, where we assumed there are no singularities. In the lower-half plane, there are always two poles from the external retarded graviton propagator.
@@ 4.2.2 Coulomb region: subleading-soft and peeling violation
We argued that, because gravitons in the Coulomb region have momenta $|\vec{x}|^{-1}$, their contribution to the metric is governed by the soft expansion of the source. Having discussed the leading-order contributions in the previous section, we now proceed to discuss the first subleading contributions. We will find agreement with the results of Refs. Christodoulou:2002 .
Similarly to the leading soft terms, the form of the subleading soft terms are universal Saha:2019tub , independent of the details of the scattering process, when expressed in terms of the incoming and outgoing momenta: [^21] The form of the subleading soft theorem may not look covariant. On the other hand, all naively covariant deformations are all equivalent to this order in the soft expansion, e.g. $\log k\cdot Q\sim\log\omega+\mathcal{O}(\omega^{0})$., [^22] In the following equation, the sums run from the one to five, where fifth element is meant to be the sum over emitted radiation. Whenever the emitted radiation is appearing more then once, the integration weigted by the spectrum is to supposed to be overall, as in equation (4.1).
@@
and $Q_{3}^{\prime\mu}=K^{\mu}$ is the total radiated energy during the scattering and the dots stand for the leading soft contributions. These terms originate both from the off-shell momentum-space source as well as from the Fourier-transform to impact-parameter space, from the ultra-soft-exchange region. [^23] On the last line we introduced $\ln(\omega+i\epsilon)\rightarrow\frac{1}{2}\ln(k^{2}+i\epsilon)$ because of the proportionality with the leading order soft factor and because of the consistency of the coefficients with Weinberg’s exponentiation of IR divergences. We also note that the $i\epsilon$ deformations of singularities are such that the integration contour is not pinched between these logarithmic branch points and the poles of the external propagator.
The soft theorem in Eq. (75) has been derived in $d=4$. This is a clear obstruction to performing this computation within dimensional regularization and the method of regions. Thus, in this case, we are forced to compute the contribution of the Coulomb soft region using cut-off regularization. Power counting suggests that such integrals are UV power-divergent as the cutoff is taken $\lambda\rightarrow+\infty$. As discussed below Eq. (33), such divergences should cancel when the Coulomb and radiation regions are added up, leaving finite terms that depend on the parameters of the integral. Furthermore, we will ignore the retarded contributions (i.e. with denominators $(Q_{i}\cdot k+i\epsilon)$ and $\log(\omega+i\epsilon)$) because they vanish in dimensional regularization, as discussed around Eq. (74). Thus, in other regularizations, we expect them to be UV sensitive and cancel completely against counterparts in the radiation region.
@@
for the Weyl tensor, where the restriction to $t=|\vec{x}|$ is a consequence of expanding in the Coulomb region. We notice that, by computing the $\omega$ integration by Cauchy’s theorem, we are sensitive only to the on-shell poles of the retarded graviton propagator because $t=|\vec{x}|>0$. Thus, all off-shell terms ($\propto k^{2}$) give vanishing contributions. One can show via direct evaluation in dimensional regularization that all the integrals with retarded matter propagators (i.e. with denominators $(Q_{i}\cdot k+i\epsilon)$ and $\log(\omega+i\epsilon)$) give vanishing integrals, see Eq. (74). [^26] As already briefly mentioned below Eq. (76), this statement hides some subtleties. Indeed, by using a different regulator than dimensional regularization, one may find that the retarded contributions give a non-vanishing contribution. Such terms are, however, expected to cancel exactly with analogous regulator-dependent terms in the radiation region. Thus, the details of the outgoing particles – in particular the impulse – do not contribute to the Coulomb region, recovering the conclusion of the scaling argument that the Coulomb (and U-SE) region is tree-level exact, from a different angle.
@@
As briefly mentioned in footnote 24, to enforce integration in the Coulomb region, we relate the cutoff $\lambda$ and $|\vec{x}|$ as $\lambda=\frac{\Lambda}{|\vec{x}|}$ with $\Lambda\gg 1$. The resulting metric is polyhomogeneous, of the form discussed in e.g. Ref. Geiller:2024ryw , with an overall power of $\frac{1}{|\vec{x}|^{2}}$ for the metric and $\frac{1}{|\vec{x}|^{4}}$ for the linearized Weyl tensor, with $\log|\vec{x}|$ enhancements.
To understand which terms may cancel between Eqs. (84) and (84) and their counterparts for $\mathsf{k}>\lambda$, we explore how coefficients of the various tensor structures change in these equations under rescalings and shifts of the cutoff $\lambda$. To this end, we focus on the vector integral contributing to the metric, Eq. (84). Using the relation explained in footnote 24 between $\lambda$ and $|\vec{x}|$, $\lambda=\frac{\Lambda}{|\vec{x}|}$ with $\Lambda\gg 1$ that enforces the restriction to the Coulomb region, it is easy to see that only terms proportional to $L^{\mu}$ can be changed by $\Lambda\rightarrow a\Lambda+b$ with fixed $a$ and $b$. Thus, all terms proportional to $N^{\mu}$ are scheme-independent and should not be expected to be affected by radiation-region contributions that cancel the cutoff dependence. These are the terms contributing to $\Psi_{0}$ and $\Psi_{1}$ NP scalars.
A similar analysis of the two-tensor integral in Eq. (85) contributing to the Weyl tensor suggests a much stronger scheme dependence. However, since the linearized Weyl tensor is simply given by suitably antisymmetrized derivatives of the metric, its scheme dependence is substantially weaker than the tensor-integral analysis suggests. In particular, all terms that contain an $N^{\mu}$ coming from the metric are scheme-independent. The $N^{\mu}N^{\nu}$ structure has this property.
@@
the peeling violation of Ref. Damour:1985cm depends on the $A_{ij}$ and that of Ref. Christodoulou:2002 depends on $B_{ij}$. We expect that the result of Ref. Damour:1985cm also originates from external gravitons in the Coulomb region. In particular, a dimensional analysis argument paralleling the one outlined at the beginning of this section suggests that a further source of peeling violation could be the subleading term in the soft-exchange region ($\sim\mathsf{k}^{0}$), which we have not considered in our analysis. An alternative origin of the peeling violating terms of Ref. Damour:1985cm could be terms in the Weyl tensor that are nonlinear in the metric fluctuations, in particular terms that depend on the linearized Schwarzschild metric of both incoming particles. We leave these for a future investigation.
@@
Before proceeding to the next section, let us briefly return to the scheme dependence of the contributions to $\Psi_{0}$ and $\Psi_{1}$ computed above. As we already explained, all contributions to the NP scalars coming from the terms proportional to the $N^{\mu}$ in the metric cannot be changed by adjusting the cutoff separating the Coulomb and radiation regions. The disappearance of the cutoff in Eq. (87) reinforces this conclusion. A further indication that $\Psi_{0}$, as computed above, is scheme-independent will be given in the following section, where we will show that the radiation region cannot yield $\Psi_{0}\sim 1/|\vec{x}|^{4}$ at the same order in $G$ as the Coulomb region. $\Psi_{1}$, on the other hand, exhibits explicit cutoff dependence of the schematic form $\Psi_{1}\sim G^{2}/|\vec{x}|^{4}\ln\lambda$. We expect that this dependence will cancel against divergences in the soft integration of the radiation region arising from logarithmic dependence on the retarded time. Moreover, while we kept non-logarithmic terms in Eq. (89), they should not be taken too seriously because they can be changed by altering the choice of cutoff in footnote 24.
@@ 4.3 Radiation regions: the analytic contributions
Having analyzed the contribution of the Coulomb region to the NP scalars and recovered the departure from Sachs’s peeling property found from different considerations in Refs. Damour:1985cm ; Christodoulou:2002 , we now proceed to discuss the contribution of the radiation region. We will show that, in this region, the leading $1/|\vec{x}|$ behavior of the Newman–Penrose scalars is determined entirely by the on-shell metric. Moreover, the components of the momentum-space metric that are analytic for vanishing graviton virtuality, $k^{2}=0$, yield position-space contributions that satisfy Sachs’s peeling property. The non-analytic terms will be examined in the next section.
As we discussed, the metric in Fourier space is given by the graviton form factor, with the subtraction of certain discontinuities. The external matter lines are on-shell and amputated through LSZ reduction, while the graviton line is off-shell and not amputated. Since we assume the current $J^{\mu\nu}$ is analytic at $k^{2}=0$ for the purpose of this section, the integral over $\omega$ can be evaluated using residue theorem. There are contributions from the pole of the external graviton propagator, as well as from possible singularities in $\omega$ within the form factor. The LSZ reduction guarantees that such singularities are at $k^{2}\neq 0$. We will consider them separately and find that the pole of the external propagator contributes to the leading $1/|\vec{x}|$ falloff, while all the others lead to a faster (exponential) decay.
@@ 4.3.1 The Fourier transform: retarded graviton poles
We begin by computing the contribution from the pole of the external propagator and we use the analyticity of the source to compute the $\omega$ integral; this effectively sets the source on-shell. Given Eq. (3.1), the result is
@@
Since we are in the radiation region ($k^{\mu}\sim|\vec{x}|^{0}$), we expand ${\widetilde{h}}^{\mu\nu}(k)$, or equivalently, the source $J^{\mu\nu}(k)$, in the angular variable $y$, around y=0y=0.272727The Watson’s lemma guarantees that it is sufficient to consider the asymptotic expansion of the source around such point. A rescaling of $y$ turns this into a series expansion in $|\vec{x}|^{-1}$: each $y$ numerator factor pulls down a factor of $(\mathsf{k}|\vec{x}|)^{-1}$. The expansion in $y$ also generates a polynomial in $n_{\perp}$; the integral over these variables is carried out via identities like Eq. (36). We are left with a non-trivial integral over $\mathsf{k}$.
@@
We can also compute subleading contributions to the metric in the $|\vec{x}|^{-1}$ expansion by keeping higher-order terms in the expansion in $y$. In the strict on-shell limit enforced by localizing onto the external propagator pole, the momentum-space metric $\tilde{h}^{\mu\nu}(k)$ becomes independent of $k^{2}$. This suggests that we should not expect any departure from the leading-order behavior described above, at least if the initial state contains only two matter particles and no gravitational radiation.
@@ 4.3.2 The Fourier transform: contribution of other poles
Having understood the first and last terms in Eq. (90) and their relation to the KMOC representation of the waveform, let us proceed to the second and third terms in that equation. They depend on the advanced time, which is parametrically large in the limit we are interested in, $v=t\,+|\vec{x}|\sim 2|\vec{x}|\rightarrow\infty$. Thus the $\mathsf{k}$ integral is dominated by $\mathsf{k}\sim|\vec{x}|^{-1}$ which, together with the on-shell condition imposed by the external graviton pole, implies that
$k^{\mu}\sim|\vec{x}|^{-1}\ .$ (92)
Thus, these integrals are dominated by graviton momenta in the Coulomb region, which we already discussed in Secs. 4.2.1 and 4.2.2. In the radiation region, where $k^{\mu}\sim|\vec{x}|^{0}$, their contribution is exponentially suppressed by the Fourier phase factor.
We now discuss the contribution of other possible poles in $\omega$ that originate from the momentum-space metric ${\widetilde{h}}^{\mu\nu}(k)$. All such poles will set $\mathsf{k}=a\omega$ where $a$ is some function of scalar products. This coefficient cannot be unity because because such a value also sets the external graviton on-shell and such a higher-order pole is in disagreement with known properties of 5-point S-matrix elements. The residue of the Fourier-transform integrand at this pole is proportional to $\exp[\pm i\omega(t\pm a|\vec{x}|)]$. The exponent is parametrically large in all terms. Thus, as for the advanced poles, the contribution of these poles to the Fourier transform is from the Coulomb region, $k^{\mu}\sim|\vec{x}|^{-1}$. It is exponentially suppressed in the radiation region.